OpenStudy (anonymous):

What is the solution set of the equation x + 2 = 4/x -2?

OpenStudy (anonymous):

\[ x+2=\frac 4{x-2} \]

OpenStudy (anonymous):

You can multiply both sides by \(x-2\): \[ x\neq 2, (x+2)(x-2) = 4\\ x^2-4=4\\ x^2=8 \]